Hi, experts/lifesavers...Can you please assist
THE SUM OF TWO NUMBERS IS 108. THE SECOND IS 8 MORE THAN 4 TIMES THE FIRST. WHAT ARE THE TWO NUMBERS? THANKS

Answer by Scriptor(36)

Hi,
Call the first number x and the second number y.
We know that:
x + y = 108
and we also know that:
y = 4x + 8 (the second is 8 more then 4 times the first)
Now fill in that second y in the first line:
x + (4x + 8) = 108
5x + 8 = 108
5x = 100
x = 20
=> y = 4*20 + 8 = 88
Check for yourself that these 2 numbers are the right ones.